to win a lottery you have to correctly pick the set of 6 numbers drawn from numbers 1-52 what is the probability of winning if you buy one ticket
Answer #1

6 numbers can be drawn from 52 numbers in C(52, 6) ways = 20358520 ways

Only one set of 6 numbers will match the drawn numbers and win the lottery

So, probability = 1/20358520 = 4.9119 * 10^-8

[Please give me a Thumbs Up if you are satisfied with my answer. If you are not, please comment on it, so I can edit the answer. Thanks.]